OpenStudy (anonymous):

7/5N + 3/5= -N HOW TO WORK

OpenStudy (amistre64):

multiply by 5, subtract by 3, add by 5N, combine like terms, and divide off the 12 perhaps

OpenStudy (amistre64):

\[\frac{7}{5}N+\frac{3}{5}=-N\] \[\frac{7(5)}{5}N+\frac{3(5)}{5}=-(5)N\] \[7N+3-3=-5N-3\] \[7N+5N=-3\] \[12N=-3\] \[\frac{12N}{12}=\frac{-3}{12}\] \[N=-\frac{3}{12}=-\frac{1}{4}\]

OpenStudy (amistre64):

algebra is alot like shuffling cards

OpenStudy (anonymous):

why didnt the 3(5) make it 15?

OpenStudy (amistre64):

becasue 5/5 = 1 in the end of it. but dont let that confuse you; we could have gone thru that step just as well. \[\frac{3(5)}{5}=\frac{15}{5}=3\]

OpenStudy (anonymous):

ok i am still skipping steps

OpenStudy (amistre64):

just take practice :)